OpenStudy (anonymous):

A waiter earned $16 in tips in an hour. He was paid in $1 bills and $5 bills. Use a model to determine how many $5 bills he received if he received a total of 12 bills. 1 2 3 4

OpenStudy (anonymous):

let x = number of 1$ bills let y = number of 5$ bills x+y=12 1*x+5*y=16 solve for y?

OpenStudy (anonymous):

$16 in $1 bills and $5 bills." x = number of one-dollar bills y = number of five-dollar bills 1x + 5y = 16 "he received a total of 12 bills" x + y = 12 x = 12 - y 1(12-y) + 5y = 16 12 - y + 5y = 16 4y = 4 y = 1 x = 12 - y = 11

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

whats your answer options?

OpenStudy (anonymous):

1 2 3 4

OpenStudy (anonymous):

ohhh sorry idk!!!! google it i did for ya!

OpenStudy (anonymous):

its asking for number of 5 dollar bills, you've determined there are 11 1$ bills out of 12 bills so that must mean there are ___ 5$ bills?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

try it! will it tell you if your right or wrong?

OpenStudy (anonymous):

yup it said 1

OpenStudy (anonymous):

so its right?